An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation of $78.11. (a) At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket? State your hypotheses and decision rule. (b) Is this a close decision?
Answer by stanbon(48568)

An auditor reviewed 25 oral surgery insurance claims from a particular surgical office, determining that the mean out-of-pocket patient billing above the reimbursed amount was $275.66 with a standard deviation of $78.11.
(a) At the 5 percent level of significance, does this sample prove a violation of the guideline that the average patient should pay no more than $250 out-of-pocket? State your hypotheses and decision rule.
---
Ho: u = 250
Ha: u > 250
----
Critical value for a one-tail test with alpha = 5%: z = 1.645
---
Test statistic: z(275.66) = (275.66-250)/[78.11/Sqrt(25)] = 1.6456
-------------------
Conclusion: Since the TS is not in the reject interval, Fail to reject Ho.
-------------------
(b) Is this a close decision?
Yes, the test statistic is very close to the critical value.
p-value = P(z > 1.6456) = 0.05024, which is very close to the alpha value.
There are only 5.024% of test results that could have provided stronger
evidence for rejecting Ho.
============================
Cheers,
Stan H.